# An EFT for anisotropic anti-ferromagnets: gapped Goldstones, pseudo-Goldstones, and phase transitions

### v1

@@ II.2 EFT with anisotropies
As anticipated, we are interested in extending the EFT presented above to include the effects of tiny explicit breaking of the internal SO(3) symmetry, which are at the origin of the observed magnon gaps (e.g., Hutchings and Samuelsen, 1971, 1972). As mentioned in the Introduction, we are interested in the case of “easy plane” anti-ferromagnets, which involve two distinct anisotropies. (For a simpler illustrative case, see Brauner (2024).) In particular, the microscopic Hamiltonian is modified to Keffer and Kittel (1952); Hutchings and Samuelsen (1972),
$H=H_{B}+\sum_{\bm{i}}D_{x}\left(\bm{\mathcal{S}}_{\bm{i}}^{x}\right)^{2}-\sum_{\bm{i}}D_{z}\left(\bm{\mathcal{S}}_{\bm{i}}^{z}\right)^{2}\,,$ (10)
where $D_{x}\,,D_{z}>0$ are new anisotropic energies. The anisotropies along the $x$- and $z$-axes typically go under the name of, respectively, “hard axis” and “easy axis” anisotropy Rezende et al. (2019).
@@
$\begin{split}\mathcal{L}={}&\frac{c_{1}}{2}\Big{[}{(\partial_{t}\hat{\bm{n}}+\mu\bm{B}\times\hat{\bm{n}})}^{2}-{v_{\theta}^{2}}{(\nabla_{i}\hat{\bm{n}})}^{2}\\ &+2{\lambda_{z}}\hat{n}_{z}^{2}-2{\lambda_{x}}\hat{n}_{x}^{2}\Big{]}\,,\end{split}$ (11)
where $\lambda_{x}\,,\lambda_{z}>0$ are two new effective coefficients. Their sign has been chosen to reproduce those in Eq. (10).
@@
$\mathcal{H}\big{|}_{\rm{stat.},\,\rm{homog.}}={}$ $-\frac{c_{1}}{2}\Big{[}(\mu\bm{B}\times\hat{\bm{n}})^{2}+2\lambda_{z}\hat{n}_{z}^{2}-2\lambda_{x}\hat{n}_{x}^{2}\Big{]}$
@@
Now, since $\lambda_{x}>0$, the second term in the expression above is necessarily minimized by setting $\hat{n}_{x}=0$. For the first term, instead, we have two possible configurations, depending on the relative strength between magnetic field effects and anisotropic ones. In particular, by defining a “spin flop” field, $B_{\rm s.f.}\equiv\sqrt{2\lambda_{z}}/\mu$, we have,
  • $\circ$
    for $B>B_{\rm{s.f.}}$ the first terms is minimized by $\hat{n}_{z}=0$, and the ground state is
    $\langle\hat{\bm{n}}\rangle=\hat{\bm{y}}\,;$
  • $\circ$
    for $B<B_{\rm{s.f.}}$ the first term is minimized by $\hat{n}_{z}=1$, and the ground state is
    $\langle\hat{\bm{n}}\rangle=\hat{\bm{z}}\,.$
This situation is schematically represented in Figure 1. Let us study these two phases separately, starting with the instance where the anisotropies are a small correction to what done so far, and then consider the case where they become dominant over the magnetic field.
[…]
Figure 1: Background configuration in the simultaneous presence of an external magnetic field $\bm{B}=B\hat{\bm{z}}$, and both easy and hard axis anisotropies. Depending on the magnitude of the external magnetic field $B$, the system undergoes a phase transition, which, in the EFT framework, consists in a discontinuous change of the direction of the ground state.
II.3 Small anisotropies: the $B>B_{\rm s.f.}$ phase
In this case the background is the same as in the absence of anisotropies, and we thus expect the explicit breaking effect to simply contribute to a small gap to the Goldstone modes. Indeed, by expanding again at quadratic order in the fluctuations, one gets,
$\begin{split}\mathcal{L}={}&\frac{c_{1}}{2}\bigg{[}{(\dot{\theta}^{1})}^{2}-v_{\theta}^{2}{(\bm{\nabla}\theta^{1})}^{2}-(\mu^{2}B^{2}-2\lambda_{z}){(\theta^{1})}^{2}\\ &+{(\dot{\theta}^{3})}^{2}-v_{\theta}^{2}{(\bm{\nabla}\theta^{3})}^{2}-2\lambda_{x}{(\theta^{3})}^{2}+\mathcal{O}\big{(}\theta^{4}\big{)}\bigg{]}\,.\end{split}$ (13)
@@
$\omega_{q,-}={}$ $\sqrt{\mu^{2}B^{2}-2\lambda_{z}+v_{\theta}^{2}q^{2}}\,.$ (14b)
As anticipated, in this phase, the role of the anisotropies is that of turning the exact Goldstones described in Section II.1 into pseudo-Goldstones.
II.4 Large anisotropies: the $B<B_{\rm s.f.}$ phase
When the effects of the anisotropies are larger than those of the applied magnetic field, the background changes, and the system undergoes a phase transition. As an indication of this, the gap of the mode in Eq. (14b) becomes imaginary, thus pointing to an instability. As anticipated, the new background is now $\langle\hat{\bm{n}}\rangle=\hat{\bm{z}}$. In this case, the broken generators are $S_{1}$ and $S_{2}$, and we must parametrize the magnon fields as $\hat{\bm{n}}(x)=e^{i\theta^{a}(x)S_{a}}\cdot\hat{\bm{z}}$, with $a=1,2$. Expanding Eq. (11) at quadratic order, one obtains,
$\begin{split}\mathcal{L}={}&\frac{c_{1}}{2}\bigg{[}\big{(}\dot{\theta}^{a}-\mu B\epsilon^{ab}\theta^{b}\big{)}^{2}-v_{\theta}^{2}\big{(}\bm{\nabla}\theta^{a}\big{)}^{2}\\ &-2\lambda_{z}\big{(}\theta^{a}\big{)}^{2}-2\lambda_{x}\delta^{a2}\delta^{b2}\theta^{a}\theta^{b}+\mathcal{O}\big{(}\theta^{4}\big{)}\bigg{]}\,.\end{split}$ (15)
@@
$\begin{split}\mathcal{M}^{ab}(q,\omega)={}&c_{1}\bigg{[}\delta^{ab}\left(v_{\theta}^{2}{q}^{2}-\omega^{2}+{2\lambda_{z}}-\mu^{2}B^{2}\right)\\ &+2\lambda_{x}\delta^{a2}\,\delta^{b2}+2\,i\mu B\,\omega\,\epsilon^{ab}\bigg{]}\,.\end{split}$ (16)
The dispersion relations for the magnon modes are found by requiring that these equations of motion admit non-trivial solutions, that is to say, demanding that the determinant of the matrix above vanishes. After doing that, one obtains the following modes,
$\begin{split}\omega^{2}_{q,\alpha=\pm}={}&\mu^{2}B^{2}+{\lambda_{x}+2\lambda_{z}}+v_{\theta}^{2}q^{2}\\ &\pm\sqrt{\lambda_{x}^{2}+4\mu^{2}B^{2}\left({\lambda_{x}+2\lambda_{z}}+v_{\theta}^{2}q^{2}\right)}\,.\end{split}$ (17)
The corresponding gaps are,
$\begin{split}\omega^{2}_{0,\pm}={}&\mu^{2}B^{2}+{\lambda_{x}+2\lambda_{z}}\\ &\pm\sqrt{\lambda_{x}^{2}+4\mu^{2}B^{2}\left({\lambda_{x}+2\lambda_{z}}\right)}\,,\end{split}$ (18)
which, for zero external field, are,
$\omega_{0,+}\big{|}_{B=0}={}$ $\sqrt{2(\lambda_{x}+\lambda_{z})}\,,$ (19a)
$\omega_{0,-}\big{|}_{B=0}={}$ $\sqrt{2\lambda_{z}}\,.$ (19b)
We can understand this spectrum using the following argument. For $B=0$ and neglecting the easy-axis anisotropy (indeed, as we will see, for NiO crystals $\lambda_{z}\ll\lambda_{x}$), the ground state must have $\hat{n}_{x}=0$, i.e., performing rotations around the $x$-axis costs zero energy and, to this residual SO(2) invariance, there will be associated an exact Goldstone boson. However, introducing even a small anisotropy along the $z$-axis breaks this residual invariance, forcing the ground state along the $z$-direction. [^5] The high energy physicist would call this “vacuum selection”. We notice that the dispersion relations obtained in the isotropic case (8a) cannot be obtained taking the limit $\lambda_{x,z}\to 0$ in (17), since, in this limit, the background changes and the field parametrization must be re-evaluated.
@@ III Quantization of the EFT
We are now ready to proceed with the quantization of our theory in the non-trivial case of $B<B_{\rm s.f.}$. The Lagrangian (15) contains both a term with a single time derivative and one with two time derivatives. This makes it impossible to diagonalize the quadratic action by a local field redefinition. (See Hui et al. (2024) for a recent discussion.) This can be seen, for example, from the fact that the eigenvectors of the matrix in Eq. (16) present non-analyticities in the frequency $\omega$. In order to deal with local fields, we then continue to work in the non-diagonal basis. In this basis, the magnon fields, $\theta^{a}$, are not in one-to-one correspondence with the physical magnon states, $|\bm{q},\alpha\rangle$, which are (approximately) asymptotic states with dispersion relations $\omega_{q,\alpha}$, given in Eq. (17). This makes the quantization of the theory in the $B<B_{\rm s.f.}$ phase non-trivial.
In the following, we will denote with $a_{\bm{q},\alpha}$ the annihilation operator associated with the single magnon state |$\ket{\bm{q},\alpha}$, that is to say that,
$a_{\bm{q},\alpha}^{\dagger}\ket{0}=\ket{\alpha,\bm{q}}\,,$ (20)
where the vacuum state is the one defined by the background configuration in presence of a magnetic field, as discussed in Section II.2. We also adopt the relativistic normalization of the one-particle states, meaning the following commutation relation,
@@
$\begin{split}\theta^{a}(x)={}&\sum_{\alpha=\pm}\int\frac{d^{3}q}{(2\pi)^{3}2\omega_{q,\alpha}}e^{-i\omega_{q,\alpha}t+i\bm{q}\cdot\bm{x}}\,\mathcal{Z}^{a}_{q,\alpha}\,a_{\bm{q},\alpha}\\ &+\rm{h.c.}\,,\end{split}$ (22)
where we introduced the overlap functions (e.g., Esposito et al., 2020; Hui et al., 2024; Cheung et al., 2022; Creminelli et al., 2024), $\mathcal{Z}_{q,\alpha}^{a}$, which connect the $\theta^{a}$ fields to the physical magnon state, |$\ket{\bm{q},\alpha}$. They are defined as,
$\braket{0}{\theta^{a}(x)}{\bm{q},\alpha}=e^{-i\omega_{{q},\alpha}t+i\bm{q}\cdot\bm{x}}\,\mathcal{Z}^{a}_{q,\alpha}\,.$ (23)
Since our theory is invariant under spatial rotations but not under boosts Creminelli et al. (2024), the overlap functions will depend on $q$. To determine them, we impose the equal time commutators between the magnon fields and their conjugate momenta,
@@
where $\pi^{a}\equiv\dot{\theta}^{a}-\mu B\,\epsilon^{ab}\,\theta^{b}$. From the above commutators we get, respectively, the conditions,
$\sum_{\alpha}\frac{\mathcal{Z}^{a}_{q,\alpha}\mathcal{Z}^{b\,*}_{q,\alpha}-\mathcal{Z}^{a*}_{q,\alpha}\mathcal{Z}^{b}_{q,\alpha}}{\omega_{q,\alpha}}$ $=0\,,$ (25a)
$\sum_{\alpha}\left[\mathcal{Z}^{a}_{q,\alpha}\mathcal{Z}^{b*}_{q,\alpha}+\mathcal{Z}^{a*}_{q,\alpha}\mathcal{Z}^{b}_{q,\alpha}\right]$ $=2\delta^{ab}\,,$ (25b)
$\sum_{\alpha}{\omega_{q,\alpha}}\left[\mathcal{Z}^{a}_{q,\alpha}\mathcal{Z}^{b*}_{q,\alpha}-\mathcal{Z}^{a*}_{q,\alpha}\mathcal{Z}^{b}_{q,\alpha}\right]$ $=4i\mu B\epsilon^{ab}\,.$ (25c)
@@
We can use the equation above to solve for one component of the overlap functions in terms of the other. In particular, we get,
$\mathcal{Z}_{q,\pm}^{2}={}$ $i\left(\frac{v_{\theta}^{2}q^{2}-\mu^{2}B^{2}+2\lambda_{z}-\omega_{q,\pm}^{2}}{2\mu B\>\omega_{q,\pm}}\right)\mathcal{Z}_{q,\pm}^{1}\,.$ (27)
Plugging this into the constraints reported in Eqs. (25) one finally finds,
${|\mathcal{Z}^{1}_{q,\pm}|}={}$ $\sqrt{\frac{1}{2}\pm\frac{2\mu^{2}B^{2}-\lambda_{x}}{2\sqrt{4\mu^{2}B^{2}(v_{\theta}^{2}q^{2}+\lambda_{x}-2\lambda_{z})+\lambda_{x}^{2}}}}\,.$ (28)
@@
We notice that, taking first the limit $B\rightarrow 0$, and then $\lambda_{z}\rightarrow 0$, we get,
@@
To make sure of the correctness of our results, in Appendix B we re-derive the overlap functions using so-called “polology” arguments.

### v2

@@ II.2 EFT with anisotropies
As anticipated, we are interested in extending the EFT presented above to include the effects of tiny explicit breaking of the internal SO(3) symmetry, which are at the origin of the observed magnon gaps [e.g., 27, 28]. As mentioned in the Introduction, we are interested in the case of “easy plane” antiferromagnets, which involve two distinct anisotropies. (For a simpler illustrative case, see Ref. [24].) In particular, the microscopic Hamiltonian is modified to [29, 28]
$H=H_{\rm H}+\sum_{\bm{i}}D_{x}\left(\bm{\mathcal{S}}_{\bm{i}}^{x}\right)^{2}-\sum_{\bm{i}}D_{z}\left(\bm{\mathcal{S}}_{\bm{i}}^{z}\right)^{2}\,$, (10)
where $D_{x}\,,D_{z}>0$ are new anisotropic energies. The anisotropies along the $x$- and $z$-axes typically go under the name of, respectively, “hard axis” and “easy axis” anisotropy [11].
@@
[…]
where $\lambda_{x}\,,\lambda_{z}>0$ are two new effective coefficients, which can be determined in terms of the microscopic anisotropic energies (see Sec. IV). Their sign has been chosen for later convenience. In particular, the last two terms indeed explicitly break the full SO(3) group.
@@
$\mathcal{H}\big{|}_{\rm{stat.},\,\rm{homog.}}={}$ $-\frac{c_{1}}{2}\Big{[}(\mu\text{\bf H}\times\hat{\bm{n}})^{2}+2\lambda_{z}\hat{n}_{z}^{2}-2\lambda_{x}\hat{n}_{x}^{2}\Big{]}$
$={}$ $\frac{c_{1}}{2}\Big{[}(\mu^{2}\text{H}^{2}-2\lambda_{z})\hat{n}_{z}^{2}+2\lambda_{x}\hat{n}_{x}^{2}\Big{]}\,$, (12)
where we have dropped any irrelevant constant terms. Now, since $\lambda_{x}>0$, the second term in the expression above is necessarily minimized by setting $\hat{n}_{x}=0$. For the first term, instead, we have two possible configurations, depending on the relative strength between magnetic field effects and anisotropic ones. In particular, by defining a “spin-flop” field, $\text{H}_{\rm s.f.}\equiv\sqrt{2\lambda_{z}}/\mu$, we have
  • $\circ$
    for ${\rm H}>{\rm H}_{\rm{s.f.}}$ the first terms is minimized by $\hat{n}_{z}=0$, and the ground state is
    $\langle\hat{\bm{n}}\rangle=\hat{\bm{y}}\,$;
    In this phase, the background value of the spin density in Eq. (9) — i.e., the total magnetization — is $\langle\bm{s}\rangle=c_{1}{\rm H}\hat{\bm{z}}$. This is what is typically called the “spin-flop” phase [e.g., 11].
  • $\circ$
    for ${\rm H}<{\rm H}_{\rm{s.f.}}$ the first term is minimized by $\hat{n}_{z}=1$, and the ground state is
    $\langle\hat{\bm{n}}\rangle=\hat{\bm{z}}\,$.
    In this phase, instead, we have $\langle\bm{s}\rangle=0$, which identifies this as the standard antiferromagnetic phase.
This situation is schematically represented in Fig. 1. Importantly, when the external magnetic field is varied across its spin-flop value, the order parameter, $\langle\hat{\bm{n}}\rangle$, is discontinuous. This is the hallmark of a first order phase transition [30, 11]. Let us study these two phases separately, starting with the instance where the anisotropies are a small correction to what done so far, and then consider the case where they become dominant over the magnetic field.
@@
Figure 1: Background configuration in the simultaneous presence of an external magnetic field $\text{\bf H}={\rm H}\hat{\bm{z}}$, and both easy and hard axis anisotropies. Depending on the magnitude of the external magnetic field $\rm H$}, the system undergoes a phase transition, which, in the EFT framework, consists in a discontinuous change of the direction of the ground state.
II.2.1 Small anisotropies: the ${\rm H}>{\rm H}_{\rm s.f.}$ phase
In this case the background is the same as in the absence of anisotropies, and we thus expect the explicit breaking effect to simply contribute to a small gap to the Goldstone modes. Indeed, by expanding again at quadratic order in the magnon fields, one gets,
$\begin{split}\mathcal{L}={}&\frac{c_{1}}{2}\bigg{[}{(\dot{\theta}^{1})}^{2}-v_{\theta}^{2}{(\bm{\nabla}\theta^{1})}^{2}-(\mu^{2}{\rm H}^{2}-2\lambda_{z}){(\theta^{1})}^{2}\\ &+{(\dot{\theta}^{3})}^{2}-v_{\theta}^{2}{(\bm{\nabla}\theta^{3})}^{2}-2\lambda_{x}{(\theta^{3})}^{2}+\mathcal{O}\big{(}\theta^{4}\big{)}\bigg{]}\,.\end{split}$ (13)
@@
$\omega_{q,-}={}$ $\sqrt{\mu^{2}{\rm H}^{2}-2\lambda_{z}+v_{\theta}^{2}q^{2}}\,.$ (14b)
As anticipated, in this phase, the role of the anisotropies is that of turning the exact Goldstones described in Section II.1 into pseudo-Goldstones. Moreover, the quadratic action above is simply a free theory for two independent real scalar fields. Consequently, it can be quantized in the standard way.
II.2.2 Large anisotropies: the ${\rm H}<{\rm H}_{\rm s.f.}$ phase
When the effects of the anisotropies are larger than those of the applied magnetic field, the background changes, and the system undergoes a phase transition. As an indication of this, the gap of the mode in Eq. (14b) becomes imaginary, thus pointing to an instability. As anticipated, the new background is now $\langle\hat{\bm{n}}\rangle=\hat{\bm{z}}$. In this case, the broken generators are $S_{1}$ and $S_{2}$, and we must parametrize the magnon fields as $\hat{\bm{n}}(x)=e^{i\theta^{a}(x)S_{a}}\cdot\hat{\bm{z}}$, with $a=1,2$. Expanding Eq. (11) at quadratic order, one obtains
$\begin{split}\mathcal{L}={}&\frac{c_{1}}{2}\bigg{[}\big{(}\dot{\theta}^{a}-\mu{\rm H}\epsilon^{ab}\theta^{b}\big{)}^{2}-v_{\theta}^{2}\big{(}\bm{\nabla}\theta^{a}\big{)}^{2}\\ &-2\lambda_{z}\big{(}\theta^{a}\big{)}^{2}-2\lambda_{x}\delta^{a2}\delta^{b2}\theta^{a}\theta^{b}+\mathcal{O}\big{(}\theta^{4}\big{)}\bigg{]}\,.\end{split}$ (15)
@@
$\begin{split}\mathcal{M}^{ab}(q,\omega)={}&c_{1}\bigg{[}\delta^{ab}\left(v_{\theta}^{2}{q}^{2}-\omega^{2}+{2\lambda_{z}}-\mu^{2}{\rm H}^{2}\right)\\ &+2\lambda_{x}\delta^{a2}\,\delta^{b2}+2\,i\mu{\rm H}\,\omega\,\epsilon^{ab}\bigg{]}\,.\end{split}$ (16)
The dispersion relations for the magnon modes are found by requiring that these equations of motion admit non-trivial solutions, that is to say, demanding that the determinant of the matrix above vanishes. After doing that, one obtains the following modes:
$\begin{split}\omega^{2}_{q,\alpha=\pm}={}&\mu^{2}{\rm H}^{2}+{\lambda_{x}+2\lambda_{z}}+v_{\theta}^{2}q^{2}\\ &\pm\sqrt{\lambda_{x}^{2}+4\mu^{2}{\rm H}^{2}\left({\lambda_{x}+2\lambda_{z}}+v_{\theta}^{2}q^{2}\right)}\,.\end{split}$ (17)
The corresponding gaps are
[…]
which, for zero external field, are
$\omega_{0,+}\big{|}_{{\rm H}=0}={}$ $\sqrt{2(\lambda_{x}+\lambda_{z})}\,$, (19a)
$\omega_{0,-}\big{|}_{{\rm H}=0}={}$ $\sqrt{2\lambda_{z}}\,.$ (19b)
We can understand this spectrum using the following argument. For ${\rm H}=0$ and neglecting the easy-axis anisotropy (indeed, as we will see, for NiO crystals $\lambda_{z}\ll\lambda_{x}$), the ground state must have $\hat{n}_{x}=0$, i.e., performing rotations around the $x$-axis costs zero energy. There will thus be an exact Goldstone boson associated with this residual SO(2) invariance. However, introducing even a small anisotropy along the $z$-axis breaks this residual invariance, forcing the ground state along the $z$-direction. [^5] The high energy physicist would call this “vacuum selection”. We notice that the dispersion relations obtained in the isotropic case (8a) cannot be obtained taking the limit $\lambda_{x,z}\to 0$ in (17), since, in this limit, the background changes and the field parametrization must be re-evaluated.
@@ III Quantization of the EFT
We are now ready to proceed with the quantization of our theory in the non-trivial case of ${\rm H}<{\rm H}_{\rm s.f.}$. Lagrangian (15) contains both a term with a single time derivative and one with two time derivatives. This makes it impossible to diagonalize the quadratic action by a local field redefinition. (See Ref. [31] for a recent discussion.) This can be seen, for example, from the fact that the eigenvectors of the matrix in Eq. (16) present non-analyticities in the frequency $\omega$. In order to deal with local fields, we then continue to work in the non-diagonal basis. In this basis, the magnon fields, $\theta^{a}$, are not in one-to-one correspondence with the physical magnon states, $|\bm{q},\alpha\rangle$, which are (approximately) asymptotic states with dispersion relations $\omega_{q,\alpha}$, given in Eq. (17). [^6] We remind the reader that the index $\alpha=\pm$ runs over the two possible magnon degrees of freedom. This makes the quantization of the theory in the ${\rm H}<{\rm H}_{\rm s.f.}$ phase non-trivial.
[…]
where the vacuum state is the one defined by the background configuration in presence of a magnetic field, as discussed in Sec. II.2. We also adopt the relativistic normalization of the one-particle states, meaning the following commutation relation,
@@
Since the quadratic Lagrangian is non-diagonal, we have that $\braket{0|\theta^{a}|\bm{q},\alpha}\,\cancel{\propto}\,\delta^{a}_{\alpha}\,$. To proceed with the quantization procedure, we first write the mode expansion of our fields,
@@
[…]
$\braket{0|\theta^{a}(x)|\bm{q},\alpha}=e^{-i\omega_{{q},\alpha}t+i\bm{q}\cdot\bm{x}}\,\mathcal{Z}^{a}_{q,\alpha}\,.$ (23)
Since our theory is invariant under spatial rotations but not under boosts [34], the overlap functions will depend on $q$. To determine them, we impose the equal time commutators between the magnon fields and their conjugate momenta,
@@
where $\pi^{a}\equiv\dot{\theta}^{a}-\mu\,{\rm H}\,\epsilon^{ab}\,\theta^{b}$. From the above commutators we get, respectively, the conditions,
$\sum_{\alpha=\pm}\frac{\mathcal{Z}^{a}_{q,\alpha}\mathcal{Z}^{b\,*}_{q,\alpha}-\mathcal{Z}^{a*}_{q,\alpha}\mathcal{Z}^{b}_{q,\alpha}}{\omega_{q,\alpha}}$ $=0\,$, (25a)
$\sum_{\alpha=\pm}\left[\mathcal{Z}^{a}_{q,\alpha}\mathcal{Z}^{b*}_{q,\alpha}+\mathcal{Z}^{a*}_{q,\alpha}\mathcal{Z}^{b}_{q,\alpha}\right]$ $=2\delta^{ab}\,$, (25b)
$\sum_{\alpha=\pm}{\omega_{q,\alpha}}\left[\mathcal{Z}^{a}_{q,\alpha}\mathcal{Z}^{b*}_{q,\alpha}-\mathcal{Z}^{a*}_{q,\alpha}\mathcal{Z}^{b}_{q,\alpha}\right]$ $=4i\mu{\rm H}\epsilon^{ab}\,.$ (25c)
@@
We can use the equation above to solve for one component of the overlap functions in terms of the other. In particular, we get
$\mathcal{Z}_{q,\pm}^{2}={}$ $i\left(\frac{v_{\theta}^{2}q^{2}-\mu^{2}{\rm H}^{2}+2\lambda_{z}-\omega_{q,\pm}^{2}}{2\mu{\rm H}\>\omega_{q,\pm}}\right)\mathcal{Z}_{q,\pm}^{1}\,.$ (27)
Plugging this into the constraints reported in Eqs. (25) one finally finds
${|\mathcal{Z}^{1}_{q,\pm}|}={}$ $\sqrt{\frac{1}{2}\pm\frac{2\mu^{2}{\rm H}^{2}-\lambda_{x}}{2\sqrt{4\mu^{2}{\rm H}^{2}(v_{\theta}^{2}q^{2}+\lambda_{x}-2\lambda_{z})+\lambda_{x}^{2}}}}\,.$ (28)
@@
We notice that, taking first the limit ${\rm H}\rightarrow 0$, and then $\lambda_{z}\rightarrow 0$, we get
@@
To make sure of the correctness of our results, in Appendix C, we rederive the overlap functions using so-called “polology” arguments.